OpenStudy (anonymous):

3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

Ok first you want to use one equation and you need to revert it back to the y= form, do you understand?

OpenStudy (anonymous):

You mean make it read like \[y = z - x\]?

OpenStudy (anonymous):

Yes , so which equation are you going to use? You can choose one or the other it wouldnt matter.

OpenStudy (anonymous):

I kinda want to start with the top one; \[3x +8y = 12\]

OpenStudy (anonymous):

Ok so what steps would you need to take to isolate y?

OpenStudy (anonymous):

Move 3x to the other side. \[3x + 8y = 12 \] \[-3x = 12 - 3x \] \[8y = 12 - 3x\]?

OpenStudy (anonymous):

Now you need to completely isolate y so what would you do? Divide 8.

OpenStudy (anonymous):

\[\frac{ 8y = 12 - 3x }{ 8 } = y = 1.5 - .375x\]

OpenStudy (anonymous):

Oops sorry this site is acting weird... Yea thats what you would get

OpenStudy (anonymous):

Yeah, its being laggy today. Okay, so now what?

OpenStudy (anonymous):

@StudiousA ?

OpenStudy (anonymous):

Ok im back, soo sorry it was acting reallydumb, anyways I think that method would be too confusing so why not use the elimination method, have you learned that yet?

OpenStudy (anonymous):

I don't think so.

OpenStudy (cuanchi):

now you replace the value of y in the other equation

OpenStudy (cuanchi):

2x + 2 (1.5-0.375x)=3

OpenStudy (cuanchi):

2x + 3- 0.75x=3 2x-0.75x=3-3 1.25x=0 x=0

OpenStudy (anonymous):

So very sorry, i think my internet is acting dumb or something, anyways... The elimination method simply uses the negated value of the second equation x coefficent over the x coefficient of the first equation and you multiply those to their respective equations.... -2 (3x +8y=12) 3 (2x+2y=3) You distribute and get -6x-16y=-24 6x+6y=9 You then add those equations add get -10y=-15 Isolate y y=1.5 Now you have to input the value of y into each equation

OpenStudy (anonymous):

is this still part A? o.o

OpenStudy (anonymous):

When you input y to both equations you get 3x +8(1.5)=12 3x+12=12 3x=0 2x+2y=3 2x+2(1.5)=3 2x=0

OpenStudy (anonymous):

I'm sorry, I'm lost...

OpenStudy (anonymous):

Yes that part was part a

OpenStudy (anonymous):

Its okay so that whole thing i wrote was part a

OpenStudy (anonymous):

Now part b says Part B: Show that the equivalent system has the same solution as the original system of equations. So usingthe information you found out about 3x and 2x you simply input, as of this moment do you understand how you got the values for. 2x and 3 x?

OpenStudy (anonymous):

What part is confusing to you at this point? I can go back and explain.

OpenStudy (anonymous):

Basically the elimination thing. I haven't learned that before.

OpenStudy (anonymous):

Ok well let me see what other method i can use to explain it.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Why dont i explain what i did

OpenStudy (anonymous):

Its called the elimination method because you are trying in this case to ELIMINATE Aa value in this case x

OpenStudy (anonymous):

Okay. That makes more sense, so what did we eliminate?

OpenStudy (anonymous):

So the first equation of 3x+8y=12 has the value of 3x and the second one is 2x+2y=3 where the value of x is 2x so you want to find values that would cancel out 3x an 2x, and to do this you want one value to be negated so it can cancel out

OpenStudy (anonymous):

So, basically finding multiples? like fractions? 6?

OpenStudy (anonymous):

So lets do it part by part, if 3x were to cancel out from 2x you would need to find a common value which is 6 so 3 x -2 and 2 x 3

OpenStudy (anonymous):

Exactly! Good job! Okay now that we have established that as mentioned it needs to cancel out so you would have to negate one value in this case instead of. 3x x 2 it would be 3x x -2

OpenStudy (anonymous):

So far so good?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Ok and when doing one thing to a part of the equation you need to do it to all parts soo....

OpenStudy (anonymous):

I read slow, 1st percentile processing speed. So if I take a bit to reply, it's because I have to read it several times before it clicks. Just saying*

OpenStudy (anonymous):

Yea its fine, no problem haha

OpenStudy (anonymous):

Okay, so we do that to both sides of the = sign, right?

OpenStudy (anonymous):

so, \[4y = 4 - 6x \]??

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

:D and thats the answers for just part A, right?

OpenStudy (anonymous):

Wait a second...

OpenStudy (anonymous):

Uh oh..

OpenStudy (anonymous):

Oh i see what you did, sorry was answering to the previous thing. You need to use the original equations and multiply

OpenStudy (anonymous):

So as we said before we are trying to cancel out the x values soo -2 (3x+8y12)

OpenStudy (anonymous):

And distribute the -2

OpenStudy (anonymous):

Oh! I tried to combine them.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Sorry this internet again.... and I meant -2(3x+8y=12)

OpenStudy (anonymous):

So, -2 (3x+8y12) = -6x + -16y = -24?

OpenStudy (anonymous):

Yea goo job! now do the same thing to the second equation but by using the value of 3 because you are trying to cancel the x value

OpenStudy (anonymous):

oops haha i meant good job, not goo... human imperfection haha, we all have it

OpenStudy (anonymous):

Ok so what would you get?

OpenStudy (anonymous):

would I use 3 or -3?

OpenStudy (anonymous):

-3, right?

OpenStudy (anonymous):

-3 (2x + 2y = 3) = -6x - 6y = -9?

OpenStudy (anonymous):

Sorry this stupd wifi ugh anyways it would be 3, do you understand why?

OpenStudy (anonymous):

Not really..

OpenStudy (anonymous):

Well you are trying to eliminate the x value correct? so you need to find what will completely take it away when you add both equations together in this case its 3 because..

OpenStudy (anonymous):

When -2 is multiplied or distributed to each value of the equation 3x +8y= 12 you get -6

OpenStudy (anonymous):

So that I don't have -6 twice? Because I'm trying to eliminate it, so x = 0 when I add -6x and 6x together!

OpenStudy (anonymous):

for the x value and if you multiply the second equation of 2x +2y=3 by 3

OpenStudy (anonymous):

by 3

OpenStudy (anonymous):

Exactly because if it were -3 then you can"t cancel out the -6 it would add and become -12

OpenStudy (anonymous):

as of this point are you understanding what to do for the first step?

OpenStudy (anonymous):

so, 3(2x + 2y = 3) = 6x + 6y = 9

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

so now add the equations

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

I believe so. I'm trying to eliminate x from both equations. correct? also, after adding them , its -10y = 15

OpenStudy (anonymous):

Omg thats correct, good job! Now what you want to do is isolate y, so what would you do to get the value of y?

OpenStudy (anonymous):

Divide, right?

OpenStudy (anonymous):

exactly, so what would get? 15 divided by -10

OpenStudy (anonymous):

I dropped my pretzel from excitement haha

OpenStudy (anonymous):

\[\frac{-10y = 15}{-10}\] \[y = 1.5\]

OpenStudy (anonymous):

yes good, now what you are going to have to do is input that value in both original equations.

OpenStudy (anonymous):

It must feel good being able to teach someone something they wouldn't otherwise understand. (:

OpenStudy (anonymous):

so input the value of y which is 1.5 into both 3x+8y=12 and into 2x=2y=3

OpenStudy (anonymous):

Okay, um. \[3x + 8(1.5) = 12\]? I don't think we ever got an x-value?

OpenStudy (anonymous):

oops i meant 2x+2y=3 yea, i love helping others and when i see i actually had an impact on them, well the feeling is great

OpenStudy (anonymous):

You will get an x value but not one in which the x is isolated, 3x is treated as a whole value and so is 2x

OpenStudy (anonymous):

so first complete 3x+8(1.5)=12

OpenStudy (anonymous):

what did you get for it?

OpenStudy (anonymous):

Okay, let me see. \[3x + 8(1.5) = 12\] \[3x + -12 = 12\] and.. \[2x + 2(1.5) = 3\] \[2x + 3 = 12\]

OpenStudy (anonymous):

not exactly, do you want me to walk you through it?

OpenStudy (anonymous):

Oops! I messed up at the end, it should be \[2x + 3 = 3\]

OpenStudy (anonymous):

yea the value for the second is correct 2x+3=3

OpenStudy (anonymous):

but 8 x 1.5 isnt -12, it is... be careful while putting it in the calculator or doing mentally, whichever you are doing

OpenStudy (anonymous):

And that shouldnt be -12..

OpenStudy (anonymous):

yea, so what is it

OpenStudy (anonymous):

I typed it wrong, but calculated correctly. Lol The equations are: \[3x + 12 = 12\] \[2x + 3 = 3\]

OpenStudy (anonymous):

Correct! Good job! ok now as I mentioned before both 3x and 2x are treated as a single value so.. Now isolate 3x from the first equation

OpenStudy (anonymous):

What do you do?

OpenStudy (anonymous):

I subtract 12 from both sides, to get 3x on its own, but then its 3x = 0

OpenStudy (anonymous):

Correct!!! Now using the same concept do that to 2x in the second equation, what do you get?

OpenStudy (anonymous):

2x = 0